# Calibrating Bayesian Tension Statistics using Neural Ratio Estimators

### v1

@@ VIII Limitations
As with all simulation based inference methods, the success of the tensionnet is dependent on how well the simulations represent the true observed datasets. In some respects, the method is also limited by the need for simulations. For example, to assess the tension between supernova observations of $H_{0}$ and CMB measurements using the $R$ statistic and the tensionnet one would need to be able to simulate the observations in a consistent framework. While work is being pursued in this direction (e.g. Watts et al., 2023; Karchev et al., 2024) it is a notoriously difficult problem.
It is also currently difficult to verify the output of the tensionnet. In practice, one could run a coverage test on the recovered distribution of $\log R$ Lemos et al. (2023). However, this only tells you how self-consistent the recovered distribution is and not whether it is centred around the correct $\log R$ value. One way to test this is to take a number of simulated datasets in the predicted distribution and calculate their $\log R$ value via an independent method such as nested sampling. An alternative validation approach is to repeat the NRE training to check for stability as in fig. 8.
As demonstrated in section VI, the tensionnet is limited by the NREs ability to predict extreme values of $\log R$. Sensible choices of prior distributions can help alleviate this issue, and the validation methods discussed above can help build confidence in the predicted distribution.
@@ IX Conclusions
A number of ways to quantify tension have been proposed including eigentension, goodness of fit degradation and Suspiciousness and these can often be translated into $\sigma$s of tension where $\sigma$ is the standard deviation of a normal distribution. A Bayesian way to quantify tension is with the tension statistic $R$ which encodes our increased confidence in one experiment’s measured data given observations from another. Formerly, $R$ is the ratio of joint Bayesian evidence to the product of the individual evidences for two datasets under a common model and prior. It is symmetric, parameterisation invariant and dimensionally consistent, however, it has a non-trivial dependence on the prior. $\log R$ is typically interpreted as indicating tension if $R\ll 1$ and concordance if $R\gg 1$ or via a Jeffery’s scale, neither of which properly account for the prior dependence.
For any pair of experiments observing the same physics, any model for the data and any prior distribution, there is a distribution of in concordance $\log R$ values. Having access to this distribution allows you to calibrate out the prior dependence from the observed $R$ and robustly convert the statistic into $\sigma$s of tension or concordance. Unfortunately, for most problems, this distribution is not analytically accessible. In this paper, we have shown that it can be readily accessed with simulations of the experimental observables and neural ratio estimation.
@@
In DESI Collaboration (2024b) some tension has been seen between the SDSS and DESI datasets at $z_{\mathrm{eff}}\approx 0.7$. In practice, this could be assessed with the tensionnet in the future should a correlated likelihood function become available for calculating the observed $R$ with nested sampling.
@@
We have shown that neural ratio estimators offer a cheap and effective way to access the in concordance $\log R$ distribution needed to calibrate out the prior dependence of the $R$ statistic. While acknowledging the limitations of this method, we believe it offers a promising step towards simulation based tension quantification. We expect that the method proposed in this paper will be broadly applicable beyond cosmology in other fields where tensions appear (e.g. CDF Collaboration, 2022).

### v2

@@ VIII Limitations
As with all simulation based inference methods, the success of the tensionnet is dependent on how well the simulations represent the true observed datasets. In some respects, the method is also limited by the need for simulations. For example, to assess the tension between supernova observations of $H_{0}$ and CMB measurements using the $R$ statistic and the tensionnet one would need to be able to simulate the observations in a consistent framework. While work is being pursued in this direction (e.g. Watts2023Cosmoglobe; Karchev2024supernova) it is a notoriously difficult problem.
It is also currently difficult to verify the output of the tensionnet. In practice, one could run a coverage test on the recovered distribution of $\log R$ Lemos2023CoverageTest. However, this only tells you how self-consistent the recovered distribution is and not whether it is centred around the correct $\log R$ value. One way to test this is to take a number of simulated datasets in the predicted distribution and calculate their $\log R$ value via an independent method such as nested sampling. An alternative validation approach is to repeat the NRE training to check for stability as in fig.˜8.
As demonstrated in section˜VI, the tensionnet is limited by the NREs ability to predict extreme values of $\log R$. Sensible choices of prior distributions can help alleviate this issue, and the validation methods discussed above can help build confidence in the predicted distribution.
@@ IX Conclusions
[…]
@@
In adame2024desiTension some tension has been seen between the SDSS and DESI datasets at $z_{\mathrm{eff}}\approx 0.7$. In practice, this could be assessed with the tensionnet in the future should a correlated likelihood function become available for calculating the observed $R$ with nested sampling.
@@
We have shown that neural ratio estimators offer a cheap and effective way to access the in concordance $\log R$ distribution needed to calibrate out the prior dependence of the $R$ statistic. While acknowledging the limitations of this method, we believe it offers a promising step towards simulation based tension quantification. We expect that the method proposed in this paper will be broadly applicable beyond cosmology in other fields where tensions appear (e.g. CDF2022).